Grade 12General Physics

intensity of gravitational field where gravitational potential is constant

sreekumar
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
 

The space around a body where the gravitational force exerted by it can be experienced by any other particle is known as the gravitational field of the body. The strength of this gravitational field is referred to as intensity, and it varies from point to point.

Consider the gravitational field of a particle of mass m located at the origin (O).

Suppose that a test particle of mass m0 is placed at the point P(x, y, z). The force of gravitational attraction exerted on the test particle is given by,

        f-vectorg = (GMmo/r2r-unit

where the position vector op-vector = r,

        r = OP = |op-vector| = |r-vector| = r

and the unit vector, r-unit = r-vector / r

The intensity of this gravitational field at a point (P) is given by the force per unit mass on a test particle kept at P, i.e.

         e-vector = f-vectorg/mo

intensity-of-gravitational-field-at-point-P 

where e-vector is the gravitational intensity and  f-vectorg  is the gravitational force acting on the mass m0. The gravitational field is, therefore, a vector field.

The gravitational field at P due to a particle of mass m kept at the point O (origin) is given by

        e-vector =  f-vectorg/mo = {-(Gmmo/r2)r-unit} * 1/mo = Gm/r2  r-unit  

where r-vector = xi + yj + zk represents the position vector of the point P with respect to the source at the origin and r-unit = r-vector / r represents the unit vector along the radial direction.

The superposition principle extends to gravitational field (intensities) as well:

        e-vector = e-vector1 + e-vector2 + e-vector3 +....+ e-vectorn

wheree-vector1, e-vector2,.... e-vectorn are the gravitational field intensities at a point due to particle 1, 2, ......, n respectively.

For a continuously distributed mass, the formula changes to e-vector =∫de-vector, where de-vector  gravitational field intensity due to an elementary mass dm.